# Gaussian Splatting on the Move: Blur and Rolling Shutter Compensation for Natural Camera Motion

### v1

@@ Abstract
[Uncaptioned image]ptpInput[Uncaptioned image]ptpMotion blur[Uncaptioned image]ptpRolling shutter[Uncaptioned image]ptpOur clean renderRolling shutter effectsMotion blurMotion blurMotion blur[Uncaptioned image]VIO
@@ 1 Introduction
Most of the existing methods aiming to compensate for these effects use classical or deep learning–based approaches to recover sharpened versions of the input images, without the aid of an underlying 3D image formation model for the scene. The recent 3D novel view synthesis methods, including NeRF  [27] and 3DGS  [13], allow for an alternative approach where a sharp 3D reconstruction is recovered without manipulating the training image data as an intermediary step. Deblur-NeRF [25], and BAD-NeRF [41] perform this in the context of NeRFs and [17] in the context of 3DGS.
This work offers an alternative deblurring and rolling shutter correction approach for the 3DGS framework. Instead of learning blurring kernels from the data, as in [25], we directly model the image formation process with camera motion and rolling shutter effects leveraging velocity estimates computed using visual-inertial odometry (VIO), a technique that fuses inertial measurement unit (IMU) data with monocular video.
Figure 1: 3DGS reconstructions from the synthetic cozyroom scene under different simulated effects with (bottom row) and without (top row) our compensation. The corresponding numerical results are given in Table 1. The bottom right image is visually indistinguishable from the reference (PSNR 35.9).
We formulate an efficient differentiable motion blur and rolling-shutter capable rendering pipeline that utilizes screen space approximation to avoid recomputing or hindering the performance of 3DGS operations. To address the ill-posed nature of the de-blurring problem (cf. [11]), we leverage the regularization capabilities of the differentiable 3DGS framework and further add regularization and priors from sensor data to reduce artefacts and increase the quality of the results (see Fig. 1 for examples).
@@
Figure 2: Further 3DGS reconstructions from the synthetic factory scene. The Splatfacto method acting as a baseline. The corresponding numerical results are given in Table 1.
@@ 2 Related Work
Deblurring differentiable implicit representations for novel view synthesis is a relatively new topic and has been previously studied in [25, 5, 41] for NeRF [27] representations and [17] in the 3DGS [13] context. Both [25] and [17] incorporate additional learnable parameters that model blurring effects as part of the rendering pipeline in the form of small Multi-Layer Perceptrons (MLPs) added to the baseline models. These methods treat the training images as fixed in time and focus on extracting sharper images from blurred inputs. An alternative method is to model the per frame image formation model by accounting for the motion trajectory during the capture process as presented in [41]. Our approach is most similar to [41] in the sense that we explicitly model the blur formation process by integrating information over a short camera trajectory. Rolling shutter effects, which are not considered in the aforementioned works, is separately considered in the NeRF context in [18], which also incorporates additional learnable parameters to the training process. In contrast, our approach does not include any additional trainable parameters nor MLPs to the 3DGS pipeline, and instead uses information readily available from IMU data captured from Visual-Inertial Odometry (VIO) pipelines.
@@ 3 Methods
We start with setting notation through an overview of Gaussian Splatting as a foundation for differentiable rendering, followed by our novel method of modelling motion blur and rolling shutter effects as camera motion. For this, we also introduce screen space approximation and pixel velocity-based rasterization, concluding with the regularization strategies employed to refine our results.
@@ 3.1 Gaussian Splatting
Gaussian Splatting (3DGS, [13]) serves as an example and current cornerstone of differentiable rendering, enabling the differentiation of a pixel’s colour in relation to model parameters. This capability allows for the optimization of a loss function by comparing synthetic images to real reference images. In essence, Gaussian Splatting maps the 3D scene into 2D images through a series of Gaussian distributions, each representing a point in the scene with specific attributes such as position, shape, transparency, and colour.
@@
comparing the synthetic images $C_{i}$ to a set of $N_{\rm img}$ real reference images $C_{i}^{\prime}$. We denote by $P_{i}\in\mathrm{SE}(3)$ camera pose corresponding to image $i$ and assume $\mathcal{C}=\mathbb{R}^{3}$ for RGB colours.
@@ 3.2 Blur and Rolling Shutter as Camera Motion
where $(v_{i},\omega_{i})$ are the linear and angular velocities of the camera in its local coordinate system (assumed constant throughout the frame interval $|\Delta t|<\frac{1}{2}(T_{e}+T_{\rm ro})$).
@@ 3.3 Screen Space Approximation
In our approach, the rendering model of 3DGS is decomposed into two stages: first, transforming the Gaussian parameters from world to pixel coordinates, and then rasterizing these parameters onto the image. This transformation is pivotal for incorporating motion effects due to camera movement during frame capture. By approximating the motion in pixel coordinates, we focus on adjusting the Gaussian means to reflect this movement, simplifying the model by primarily altering these means while keeping other parameters stable. An informal sketch of this process is visualized in Fig. 3.
@@
in particular, we neglect the effect of (small) camera motion on the other view-dependent intermediary variables $(d_{i,j},\Sigma_{i,j}^{\prime},c_{i,j})$ and only model the effect on the pixel coordinates $\mu_{i,j}^{\prime}$ of the Gaussian means and introduce a new set of variables, the pixel velocities computed as
@@ 3.4 Rasterization with Pixel Velocities
Approximating camera velocity on the Gaussian means in pixel-space has the benefit that the projective transform of the 3DGS pipeline does not need to be repeated to generate each blur sample, making the method faster than treating Gaussian velocities in world space.
@@ 3.5 Pose Optimization
where $\nu$ is any component of $R_{i}$ or a parameter it depends on. This neglects the effect of small camera motions on view-dependent colours and Gaussian precisions $\Sigma_{i,j}^{\prime}$ defined in pixel coordinates, but allows approximating the expression in terms of the Gaussian position derivatives $\frac{\partial C_{i}}{\partial\mu_{j,k}}$, which are readily available in the 3DGS pipeline. This approach is similar to [45] and derived in more detail in Sec. 0.A.3.
3.6 Regularization
We investigate two additional regularization methods in the 3DGS training procedure to make our approach more robust: (i) Underestimating exposure time: $\tilde{T}_{e}=T_{e}(1-\lambda)$ (applied to synthetic and real data); (ii) Adding a different noise vector $\eta_{j}^{(k)}\sim\mathcal{\rm N}(0,(\beta|v_{j}|)^{2}I)$ to the estimated linear velocities $v_{j}^{(k)}=v_{j}+\eta_{j}^{(k)}$ on each iteration $k$ (real data only). An alternative approach to velocity regularization could be optimizing the velocity vectors as a part of 3DGS problem. This is left for future work.
@@ 4 Experiments
We evaluate the performance of our method on two different data sets: synthetic data based on the Deblur-NeRF data set [25] and real data recorded using mobile phones. We implemented our method as an extension to the open source 3DGS implementation Splatfacto in Nerfstudio [39], which is based on gsplat [47]. In particular, the implementation is based on Splatfacto in Nerfstudio version 1.0.2 and gsplat version 0.1.8 (see https://github.com/nerfstudio-project/nerfstudio/releases/tag/v1.0.2).
@@
PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$
@@
Baseline 17.11 .500 .371 15.31 .265 .430 21.17 .479 .372 14.35 .290 .467
Ours 35.92 .977 .015 29.01 .966 .019 31.57 .919 .069 31.93 .970 .014
@@ 4.1 Synthetic Data
3DGS uses a SfM-based sparse point cloud as seed points for the initialization of the Gaussian means and scales, which is typically obtained from COLMAP [32]. However, with significantly blurry and other otherwise noisy data, COLMAP often fails which prevents us from obtaining good SfM estimates for initialization on synthetic data sets like Deblur-NeRF. To overcome this, we compute an initial point cloud using SIFT [24] feature matching and triangulation with the known ground truth poses from the synthetic data set, bypassing camera pose estimation from COLMAP.
@@
We generate multiple variants of the Deblur-NeRF data set to simulate motion blur, rolling shutter, and noisy pose effects respectively. In order to obtain synthetic velocity data and simulated rolling shutter effects, we had to re-render synthethic images from the Deblur-NeRF data set [25], see Sec. 0.B.1 for further details. We experiment on the following synthetic data set variants: (i) a motion blur variant, which closely matches the motion blur variant of the original Deblur-NeRF data set, (ii) a rolling shutter only variant, (iii)  and a translational and angular pose noise variant, without RS or motion blur effects. In the motion blur and rolling shutter variants, we use the synthetic ground truth velocities $(v_{j},\omega_{j})$ without additional pose noise.
As in [25], the evaluation images in the data set are always rendered without blur or rollings shutter effects, which allows us to assess the novel-view synthesis performance of the underlying ‘sharp’ reconstruction independent of the accuracy of our simulation and compensation for these effects.
@@
We train 3DGS models for 30k iterations for all of the above synthetic data sets using 3DGS Splatfacto method as our baseline and we compare novel-view synthesis metrics with our method with the appropriate compensation mode enabled (i.e., rolling shutter compensation but no motion blur compensation or pose optimization for RS case). The results are shown in Table 1 and qualitatively visualized in Figs. 1 and 2. Our method consistently outperforms the 3DGS baseline across all scenarios, indicating the effectiveness of our approach in compensating for blurring and RS effects arising from camera motion in pixel space.
Table 2: Real data results. RS pose opt means 2-pass 3DGS pose optimization with rolling shutter compensation. This is compared to COLMAP baseline, both with (‘+ blur’) and without motion blur compensation.
COLMAP baseline COLMAP + blur RS pose opt + blur
PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$
iphone-lego1 28.04 .900 .169 28.61 .918 .128 26.40 .869 .164
iphone-lego2 27.67 .912 .133 28.38 .921 .120 24.73 .853 .195
iphone-lego3 25.16 .795 .261 25.42 .817 .210 22.93 .720 .282
iphone-pots1 28.21 .896 .184 28.29 .903 .158 27.50 .876 .169
iphone-pots2 27.16 .840 .268 27.51 .867 .203 26.32 .821 .234
pixel5-lamp 27.21 .784 .316 26.91 .783 .317 30.09 .820 .258
pixel5-plant 26.30 .902 .261 26.47 .903 .243 27.32 .908 .242
pixel5-table 28.70 .881 .310 28.88 .883 .296 31.01 .884 .278
s20-bike 26.66 .885 .270 26.68 .885 .272 28.18 .894 .262
s20-bikerack 25.71 .800 .281 25.65 .800 .284 26.68 .797 .297
s20-sign 23.42 .732 .280 24.07 .750 .272 25.89 .780 .259
average 26.75 .848 .249 26.99 .857 .228 27.00 .838 .240
@@
[…]
@@ 4.2 Smartphone Data
Manual calibration
In addition to the main sequences, we recorded separate manual camera calibration data for the Android devices, using the same fixed focus distance as in the other sequences, and used the Kalibr software package (https://github.com/ethz-asl/kalibr) to compute a accurate intrinsic calibration parameters for those devices.
@@
We first process the recorded data with the Spectacular AI SDK [1], to obtain the following: (i) A sparse set of key frames with minimum distance of 10 cm, selected to approximately minimize motion blur among the set of candidate key frames (see Sec. 0.A.5 for details). (ii) For each key frame, the frame velocities $(v_{j},\omega_{j})$, based on the fusion of IMU and video data. (iii) Approximate VISLAM-based poses $P_{j}^{\rm SAI}$.
@@
Pose estimation
After preprocessing, the poses of the key frames and camera intrinsics are estimated using COLMAP [32] (through Nerfstudio). In our main results, we only included sequences in which COLMAP did not fail due to excessively difficult visual conditions.
@@
Rolling shutter pose optimization
As an alternative pose estimation method, we use our approach with both pose optimization and rolling-shutter compensation enabled, using the factory intrinsics and rolling shutter readout time reported by the mobile devices. The reasoning behind this is that COLMAP may use the camera intrinsics to compensate also for the rolling shutter effect when used with non-still images. Similarly, COLMAP-estimated poses do not generally represent the best solution for a given (different) set of camera intrinsics, since slightly inaccurate intrinsics may be compensated by adjusting the poses. Hence enabling rolling-shutter compensation with COLMAP-estimated poses is not a good approach. We performed the experiments using both manual calibration data as well as the built-in calibration information from the devices.
Two-pass pose optimization
When optimizing the poses, it is also necessary to change adjust the poses of the evaluation images. To this end, mostly due to ease of implementation, we choose an approach where we first train the model with pose optimization using all images. We then store the adjusted poses, and then train the model again without the evaluation images, but with fixed poses. An alternative would be to train with pose optimization with training images only and then optimize the poses of the evaluation images to minimize the loss on the 3DGS model, which is not modified during this optimization.

### v2

@@ Abstract
[Uncaptioned image]pInput[Uncaptioned image]pMotion blur[Uncaptioned image]pRolling shutter[Uncaptioned image]pOur clean renderRolling shutter effectsMotion blurMotion blurMotion blur[Uncaptioned image]VIO
@@ 1 Introduction
Most of the existing methods aiming to compensate for these imaging effects use classical or deep learning–based approaches to recover sharpened versions of the input images, without the aid of an underlying 3D image formation model for the scene. The recent 3D novel view synthesis methods, including NeRF  [27] and 3DGS  [13], allow for an alternative approach where a sharp 3D reconstruction is recovered without manipulating the training image data as an intermediary step. Deblur-NeRF [25], and BAD-NeRF [42] perform this in the context of NeRFs. In the context of 3DGS, [17] propose a method for a related but different problem of de-focus blur compensation. Similarly to [42], our work focuses on camera motion blur and applications to real data where this is the most prominent blurring modality, such as hand-held captures of static scenes without extreme close-up shots.
This work offers an alternative deblurring and rolling shutter correction approach for the 3DGS framework. Instead of learning blurring kernels from the data, as in [25], we directly model the image formation process with camera motion and rolling shutter effects leveraging velocity estimates computed using visual-inertial odometry (VIO), a technique that fuses inertial measurement unit (IMU) data with monocular video. We formulate an efficient differentiable motion blur and rolling-shutter capable rendering pipeline that utilizes screen space approximation to avoid recomputing or hindering the performance of 3DGS operations. To address the ill-posed nature of the deblurring problem (cf. [11]), we leverage the regularization capabilities of the differentiable 3DGS framework priors information from sensor data.
@@
Figure 1: 3DGS reconstructions from the synthetic cozyroom scene under different simulated effects with (bottom row) and without (top row) our compensation. The corresponding numerical results are given in Table 1. The bottom right image is visually indistinguishable from the reference (PSNR 36.2).
Figure 2: Further 3DGS reconstructions from the synthetic factory scene. The Splatfacto method acting as a baseline. The corresponding numerical results are given in Table 1. In this case, the pose optimization case converged to a local minimum.
@@ 2 Related Work
Deblurring differentiable implicit representations for novel view synthesis is a relatively new topic and has been previously studied in [25, 5, 42] for NeRF [27] representations and [17] in the 3DGS [13] context. Both [25] and [17] incorporate additional learnable parameters that model blurring effects as part of the rendering pipeline in the form of small Multi-Layer Perceptrons (MLPs) added to the baseline models. These methods treat the training images as fixed in time and focus on extracting sharper images from blurred inputs. An alternative method is to model the per frame image formation model by accounting for the motion trajectory during the capture process as presented in [42]. Our approach is most similar to [42] in the sense that we explicitly model the blur formation process by integrating information over a short camera trajectory. Rolling shutter effects, which are not considered in the aforementioned works, is separately considered in the NeRF context in [18], which also incorporates additional learnable parameters to the training process. In contrast, our approach does not include additional MLPs to the 3DGS pipeline, but directly models the local camera trajectory using linear and angular velocities, for which good initial estimates are readily available from IMU data captured from Visual-Inertial Odometry (VIO) pipelines.
@@ 3 Methods
We start with a brief overview of Gaussian Splatting and establishing notation, followed by our formulation of motion blur and rolling shutter effects in the 3DGS framework. For this, we also introduce a novel screen space approximation and pixel velocity-based rasterization, concluding with pose optimization and its role in the evaluation methodology.
@@ 3.1 Gaussian Splatting
Gaussian Splatting (3DGS, [13]) serves as an example and current cornerstone of differentiable rendering, enabling the differentiation of a pixel’s colour in relation to model parameters. This capability allows for the optimization of a loss function by comparing rendered images to real reference images. In essence, Gaussian Splatting maps the 3D scene into 2D images through a set of Gaussian distributions with additional colour and transparency attributes, each contributing to the region of the scene defined by the its mean and covariance.
@@
comparing the rendered images $C_{i}$ to a set of $N_{\rm img}$ real reference images $C_{i}^{\prime}$. We denote by $P_{i}\in\mathrm{SE}(3)$ camera pose corresponding to image $i$ and assume $\mathcal{C}=\mathbb{R}^{3}$ for RGB colours.
@@ 3.2 Blur and Rolling Shutter as Camera Motion
where $(v_{i},\omega_{i})$ are the linear and angular velocities of the camera in its local coordinate system (assumed constant throughout the frame interval $|\Delta t|<\frac{1}{2}(T_{e}+T_{\rm ro})$). We add the velocities as additional optimizable parameters and set their initial values to the estimate from VIO, if available, or zero otherwise.
@@ 3.3 Screen Space Approximation
In our approach, the rendering model of 3DGS is decomposed into two stages: first, transforming the Gaussian parameters from world to pixel coordinates, and then rasterizing these parameters onto the image. This transformation is pivotal for incorporating motion effects due to camera movement during frame capture. By approximating the motion in pixel coordinates, we focus on adjusting the Gaussian means to reflect this movement, simplifying the model by primarily altering these means while keeping other parameters stable. A high-level overview of the process is presented in Fig. 3.
@@
in particular, to simplify and optimize the implementation, we neglect the effect of (small) camera motion on the other view-dependent intermediary variables $(d_{i,j},\Sigma_{i,j}^{\prime},c_{i,j})$ and only model the effect on the pixel coordinates $\mu_{i,j}^{\prime}$ of the Gaussian means and introduce a new set of variables, the pixel velocities computed as
@@ 3.4 Rasterization with Pixel Velocities
Approximating camera velocity on the Gaussian means in pixel-space has the benefit that the projective transform of the 3DGS pipeline does not need to be repeated to generate each blur sample (i.e. the terms in the outer sum), making the method faster than treating Gaussian velocities in world space.
@@ 3.5 Pose Optimization
where $\nu$ is any component of $R_{i}$ or a parameter it depends on. This neglects the effect of small camera motions on view-dependent colours and Gaussian precisions $\Sigma_{i,j}^{\prime}$ defined in pixel coordinates, but allows approximating the expression in terms of the Gaussian position derivatives $\frac{\partial C_{i}}{\partial\mu_{j,k}}$, which are readily available in the 3DGS pipeline. This approach is similar to [47] and derived in more detail in Sec. 0.A.3. To stabilize the reconstruction, we employ simple $l_{2}$ penalty terms to prevent the poses from drifting too far from their initial estimates.
3.6 Evaluation Strategy - Optimization with fixed Gaussians
Due to the well-known gauge invariance studied in classical photogrammetry (cf. [40]), the inverse rendering problem has seven undetermined degrees of freedom: rotating, scaling and translating the coordinate system in all the quantities $(\mathcal{G},\{P_{i},v_{i},\omega_{i}\}_{i})$ in consistent manner yields new solution $(\mathcal{G}^{\prime},\{P_{i}^{\prime},v_{i}^{\prime},\omega_{i}^{\prime}\}_{i})$ with equal photometric losses. While, the $l_{2}$ penalty on the poses will fix this gauge invariance, it leaves room for small perturbations: the reconstructed scene may slightly rotate and translate, which can cause the evaluation frames to be misaligned with reconstruction where the training frames and Gaussians $\mathcal{G}$ are consistent with each other.
To tackle this, we also optimize the poses and velocities of the evaluation frames with fixed gaussians $\mathcal{G}$, which can be achieved by including both test and training frames to the optimization problem, but disabling the back-propagation of gradient data to the parameters $\mathcal{G}$ (while allowing it for $P_{i},v_{i}$ and $\omega_{i}$) whenever an evaluation index $i$ is selected by the SGD optimizer.
@@ 4 Experiments
We evaluate the performance of our method on two different data sets: synthetic data based on the Deblur-NeRF data set [25] and real data recorded using mobile phones. We implemented our method as an extension to the open source 3DGS implementation Splatfacto in Nerfstudio [39], which is based on gsplat [49] (cf. App. 0.C for details).
@@
[…]
@@
Baseline 16.76 .484 .339 15.09 .251 .420 20.95 .465 .344 13.97 .271 .442
Ours 36.30 .976 .013 24.03 .745 .132 34.68 .951 .038 33.53 .979 .012
Table 2: Novel view synthesis results Deblur-NeRF data set variant in  [42]. The Tanabata scene has lower metrics due to an issue in the input data, cf. App. 0.B.
Cozyroom Factory Pool Tanabata Trolley
[…]
Splatfacto 24.93 .802 .225 21.28 .598 .440 27.88 .763 .281 18.52 .533 .433 19.47 .564 .387
MPR+Splatf. 29.26 .894 .093 23.38 .737 .246 30.96 .867 .176 22.77 .773 .227 26.49 .854 .185
Deblur-NeRF 29.88 .890 .075 26.06 .802 .211 30.94 .840 .169 22.56 .764 .229 25.78 .812 .180
BAD-NeRF 30.97 .901 .055 31.65 .904 .123 31.72 .858 .115 23.82 .831 .138 28.25 .873 .091
Ours 31.80 .945 .032 30.54 .946 .078 32.08 .890 .075 24.79 .912 .079 30.16 .933 .044
@@ 4.1 Synthetic Data
3DGS uses a SfM-based sparse point cloud as seed points for the initialization of the Gaussian means and scales, which is typically obtained from COLMAP [32]. However, with significantly blurry and other otherwise noisy data, COLMAP often fails which prevents us from obtaining good SfM point cloud for initialization on very blurry synthetic data sets like Deblur-NeRF. To overcome this, we compute an initial point cloud using SIFT [24] feature matching and triangulation with the COLMAP-estimated poses.
BAD-NeRF data set variant
To compare our performance with existing work, we use the BAD-NeRF re-render [42] of the synthetic Deblur-NeRF data set [25]. We computed the results of our method, the Splatfacto baseline and a variant where, prior to COLMAP registration, the images were deblurred with MPR [52]. Similarly to [42], we only use the blurry images for pose registration in COLMAP. The results of the experiment are presented in Table 2, where they are also compared to the previous NeRF-based methods as reported in [42], where we also note that the original Deblur-NeRF [25] has been evaluated using ground truth training poses. Our method clearly outperforms all baselines.
Re-rendered data
@@
To study the individual components of our method in more detail, we rendered new variants of the Deblur-NeRF data set to simulate motion blur, rolling shutter, and noisy pose effects respectively (see App. 0.B for details). We experiment on the following synthetic data set variants: (i) a motion blur variant, which closely matches the motion blur variant of the original Deblur-NeRF data set, (ii) a rolling shutter only variant, (iii)  and a translational and angular pose noise variant, without RS or motion blur effects. In the motion blur and rolling shutter variants, we use the synthetic ground truth velocities $(v_{j},\omega_{j})$ without additional pose noise. For these variants, we keep poses and velocities fixed during optimization. In this experiment, the ground truth poses were also used for initial point cloud triangulation instead of COLMAP-estimated poses.
@@
We train 3DGS models for all of the above synthetic data sets using 3DGS Splatfacto method as our baseline and we compare novel-view synthesis metrics with our method with the appropriate compensation mode enabled (i.e., rolling shutter compensation but no motion blur compensation or pose optimization for RS case). The results are shown in Table 1 and qualitatively visualized in Figs. 1 and 2. Our method consistently outperforms the 3DGS baseline across all scenarios, indicating the effectiveness of our approach in compensating for blurring and RS effects arising from camera motion in pixel space.
Table 3: PSNR metric ablation study and comparison for smartphone data.
[…]
iphone-lego1 28.05 28.12 29.20 28.59 28.71 29.03 23.26 29.20
iphone-lego2 27.85 27.88 27.95 27.39 28.15 28.55 26.45 27.95
iphone-lego3 23.75 23.71 24.50 24.10 24.22 23.78 22.45 24.50
iphone-pots1 28.32 28.58 29.10 28.91 28.93 29.18 24.44 29.10
iphone-pots2 27.25 27.39 28.00 27.68 26.66 27.81 23.64 28.00
pixel5-lamp 28.22 30.91 28.38 28.77 29.75 29.70 31.95 30.46
pixel5-plant 26.57 27.41 27.45 26.81 27.37 27.69 28.20 27.90
pixel5-table 28.93 30.82 29.25 30.33 31.16 31.47 32.42 31.86
s20-bike 27.35 27.74 27.57 27.58 27.58 27.72 26.62 28.93
s20-bikerack 25.98 29.39 27.92 26.23 26.09 28.92 27.77 29.74
s20-sign 23.71 25.93 24.47 25.44 25.54 26.28 24.19 26.84
average 26.91 27.99 27.62 27.44 27.65 28.19 26.49 28.59
@@
BaselineOurs (MB)Ours (MBRS)ReferenceRefer to caption1Refer to captionRefer to captionRefer to captionRefer to caption1Refer to captionRefer to captionRefer to captionRefer to caption1Refer to captionRefer to captionRefer to captionRefer to caption1Refer to captionRefer to captionRefer to caption
@@ 4.2 Smartphone Data
We first process the recorded data with the Spectacular AI SDK [1], to obtain the following: (i) A sparse set of key frames with minimum distance of 10 cm, selected to approximately minimize motion blur among the set of candidate key frames (see Sec. 0.A.5 for details). (ii) For each key frame, the initial estimates for the frame velocities $(v_{j}^{\rm VIO},\omega_{j}^{\rm VIO})$, based on the fusion of IMU and video data. (iii) Approximate VISLAM-based poses $P_{j}^{\rm SAI}$.
@@
Pose and intrinsic estimation
After preprocessing, the poses of the key frames and camera intrinsics are estimated using COLMAP [32] (through Nerfstudio). In our main results, we only included sequences in which COLMAP did not fail due to excessively difficult visual conditions. In App. 0.D we present additional experiments where manual calibration results or built-in calibration data from the devices is used in place of COLMAP intrinsics.
@@
CVR de-rolling baseline
As a baseline for rolling shutter compensation, we experimented with a variant where, prior to the COLMAP phase, the image data was de-rolled with CVR [6]. For a fair comparison, we enabled the other optimization features in our method in this test. We also note that this test was omitted with synthetic data as CVR requires access to consecutive frames, which were not available in the sparse synthetic test.
Ablation study
We also performed an ablation study where each of the main components of the method are individually disabled: (i) motion blur compensation, by setting $T_{e}=0$, $N_{\rm blur}=1$; (ii) rolling shutter compensation by setting $T_{\rm ro}=0$; (iii) pose optimization; (iv) pptimization of linear and angular velocities $(v_{i},\omega_{i})$; (v) velocity optimization initialization from VIO by setting $v_{i}=\omega_{i}=0$ instead of $v_{i}=v_{i}^{\rm VIO}$, $\omega_{i}=\omega_{i}^{\rm VIO}$.